I am trying to do my homework and do not understand the statement below.
We can use the Pythagorean Theorem to solve problems that involve right triangles. Provide an example of a day-to-day situation that involves right triangles and the use of the theorem.

Answer by stanbon(48564)

We can use the Pythagorean Theorem to solve problems that involve right triangles. Provide an example of a day-to-day situation that involves right triangles and the use of the theorem.
----------------------------
If it is 90 ft. from home plate to 1st base
and 90 ft from 1st base to 2nd base how far
does the catcher have to throw a baseball
to stop a player trying to steal 2nd base?
-----------------
EQUATION:
(throw)^2 = 90^2 + 90^2
throw = sqrt(2*90^2)
throw = 90sqrt(2)
throw is approximately 127.28 ft.
==================
Cheers,
Stan H.